# Sample size and power calculations for causal inference in observational studies

### v3

@@ A.3 Proof of Proposition 1
A.4 Proof of Proposition 2
Let $h(k)=\log\phi(k)=\log\Gamma\left(kr+\frac{1}{2}\right)-\log\Gamma(kr)+\log\Gamma\left(k(1-r)+\frac{1}{2}\right)-\log\Gamma(k(1-r))-\log k-\frac{1}{2}\log(r(1-r))$. Then, it suffices to show that $h^{\prime}(k)>0$ for all $k$ such that $12kr>$ 1 and $12k(1-r)>$ 1.
Denote by $\psi$ the digamma function, $\psi(x)=(\log\Gamma(x))^{\prime}$. Then,
$h^{\prime}(k)=r\psi\left(kr+\frac{1}{2}\right)-r\psi(kr)+(1-r)\psi\left(k(1-r)+\frac{1}{2}\right)-(1-r)\psi(k(1-r))-\frac{1}{k}.$ (A.9)
Using the equality
$\psi(x)-\psi(y)=\int_{0}^{1}\frac{t^{y-1}-t^{x-1}}{1-t}\,\mathrm{d}t,\quad\forall x,y>1,$ (A.10)
we have
$h^{\prime}(k)$ $=r\int_{0}^{1}\frac{t^{kr-1}\left(1-\sqrt{t}\right)}{1-t}\,\mathrm{d}t+(1-r)\int_{0}^{1}\frac{t^{k(1-r)-1}\left(1-\sqrt{t}\right)}{1-t}\,\mathrm{d}t-\frac{1}{k}$ (A.11)
$=\int_{0}^{1}\frac{rt^{kr-1}+(1-r)t^{k(1-r)-1}}{1+\sqrt{t}}\,\mathrm{d}t-\frac{1}{k}.$ (A.12)
Changing variable ($t=u^{2}$) yields
$h^{\prime}(k)$ $=\int_{0}^{1}\frac{2ru^{2kr}+2(1-r)u^{2k(1-r)}}{u(1+u)}\,\mathrm{d}u-\frac{1}{k}$ (A.13)
$=\frac{1}{k}\left\{\int_{0}^{1}\frac{\mathrm{d}\left(u^{2kr}+u^{2k(1-r)}\right)}{1+u}-1\right\}$ (A.14)
$=\frac{1}{k}\left\{\left.\frac{u^{2kr}+u^{2k(1-r)}}{1+u}\right|_{0}^{1}+\int_{0}^{1}\frac{u^{2kr}+u^{2k(1-r)}}{(1+u)^{2}}\,\mathrm{d}u-1\right\}$ (A.15)
$=\frac{1}{k}\int_{0}^{1}\frac{u^{2kr}+u^{2k(1-r)}}{(1+u)^{2}}\,\mathrm{d}u>0.$ (A.16)
A.5 Proof of Equation (12).
Recall $Y(z)=a_{z}W_{e}+\epsilon_{z}$ where $\epsilon_{z}\sim\mathsf{N}(\mu_{z},\sigma_{z}^{2})$. The observed quantities satisfy the following equations.
$E_{z}$ $\equiv\mathbb{E}[Y\mid Z=z]=\mathbb{E}[a_{z}W_{e}+\epsilon_{z}\mid Z=z]=a_{z}\mathbb{E}[W_{e}\mid Z=z]+\mu_{z};$ (A.17)
$S_{z}^{2}$ $\equiv\mathbb{V}[Y\mid Z=z]=\mathbb{V}[a_{z}W_{e}+\epsilon_{z}\mid Z=z]=a_{z}^{2}\mathbb{V}[W_{e}\mid Z=z]+\sigma_{z}^{2};$ (A.18)
$\rho_{z}$ $\equiv\mathrm{cor}[Y,W_{e}\mid Z=z]=\mathrm{cor}[a_{z}W_{e}+\epsilon_{z},W_{e}\mid Z=z]=\frac{a_{z}\mathbb{V}[W_{e}\mid Z=z]}{\sqrt{S_{z}\mathbb{V}[W_{e}\mid Z=z]}}.$ (A.19)
By (A.19),
$a_{z}=\rho_{z}\sqrt{\frac{S_{z}^{2}}{\mathbb{V}[W_{e}\mid Z=z]}},$ (A.20)
which, plugged into (A.17) and (A.18), yields
$\mu_{z}=E_{z}-a_{z}\mathbb{E}[W_{e}\mid Z=z],\quad\sigma_{z}^{2}=(1-\rho_{z}^{2})S_{z}^{2}.$ (A.21)
A.6 Proof of the analytical form of variance $V$ in Theorem 2.
@@ A.4 Proof of Proposition 2
$V=\mathbb{E}\left[\frac{\{Y(1)-\mathbb{E}[Y(1)]\}^{2}}{e(X)}\right]+\mathbb{E}\left[\frac{\{Y(0)-\mathbb{E}[Y(0)]\}^{2}}{1-e(X)}\right],$ (A.22)
where $e(X)=\operatorname{\mathrm{e}xpit}(W_{e})$, $Y(z)=a_{z}W_{e}+\epsilon_{z}$, $W_{e}\sim\mathsf{N}(\mu_{e},\sigma_{e}^{2})$, $\epsilon_{z}\sim\mathsf{N}(\mu_{z},\sigma_{z}^{2})$. Therefore,
@@
$\mathbb{E}\left[\frac{\{Y(1)-\mathbb{E}[Y(1)]\}^{2}}{e(X)}\right]=\mathbb{E}\left[\frac{a_{1}^{2}(W_{e}-\mu_{e})^{2}+\epsilon_{1}^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]=a_{1}^{2}\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{1}^{2}\mathbb{E}\left[\frac{1}{\operatorname{\mathrm{e}xpit}(W_{e})}\right],$ (A.23)
where
@@
$\mathbb{E}\left[\frac{1}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ $=\int\left(1+e^{-w}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(w-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}w$ (A.24)
$=\int\left(1+e^{-u-\mu_{e}}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u$ (A.25)
$=\int\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u+\int\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u+\sigma_{e}^{2})^{2}}{2\sigma_{e}^{2}}-\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)\,\mathrm{d}u$ (A.26)
$=1+\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right),$ (A.27)
and
@@
$\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ $=\int(w-\mu_{e})^{2}\left(1+e^{-w}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(w-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}w$ (A.28)
@@
$=\int u^{2}\left(1+e^{-u-\mu_{e}}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u$ (A.29)
@@
$=\int u^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u+\int u^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u+\sigma_{e}^{2})^{2}}{2\sigma_{e}^{2}}-\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)\,\mathrm{d}u$ (A.30)
@@
$=\sigma_{e}^{2}+(\sigma_{e}^{4}+\sigma_{e}^{2})\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (A.31)
@@
$\mathbb{E}\left[\frac{\{Y(1)-\mathbb{E}[Y(1)]\}^{2}}{e(X)}\right]=a_{1}^{2}\sigma_{e}^{2}+\sigma_{1}^{2}+\left[a_{1}^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{1}^{2}\right]\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (A.32)

### v4

@@ A.3 Proof of Proposition 1(i)
A.4 Proof of Proposition 1(ii)
Denote $\phi$ by $\phi(a,b)$ to indicate that $\phi$ is a function of $a$ and $b$, and factorize it as $\phi(a,b)=\tilde{\phi}(a)\tilde{\phi}(b)$, where $\tilde{\phi}(x)=\Gamma(x+0.5)/\{\sqrt{x}\Gamma(x)\}$.
To show $\phi(a,b)$ increases monotonically in $a$ (and $b$), it suffices to show that the partial derivative of $\log\phi(a,b)=\log\tilde{\phi}(a)+\log\tilde{\phi}(b)$ with respect to $a$ (and $b$) is positive. By seeing that $\log\tilde{\phi}(x)=\log\Gamma(x+0.5)-\log\Gamma(x)-\log\sqrt{x}$, we have
$\frac{\partial\log\phi(a,b)}{\partial a}=\frac{d\log\tilde{\phi}(a)}{da}=\psi(a+\frac{1}{2})-\psi(a)-\frac{1}{2a},$ (29)
where $\psi(x)=d\log\Gamma(x)/dx$ is the digamma function.
Leveraging Abramowitz and Stegun (1965) Equation (6.3.22), we have
$\psi(x)-\psi(y)=\int_{0}^{1}\frac{t^{y-1}-t^{x-1}}{1-t}\,\mathrm{d}t,\quad\forall x,y>0.$ (30)
Plugging into Equation (29) yields
$\psi(a+\frac{1}{2})-\psi(a)=\int_{0}^{1}\frac{t^{a-1}-t^{a-1/2}}{1-t}\,\mathrm{d}t=\int_{0}^{1}\frac{t^{a-1}}{1+\sqrt{t}}\,\mathrm{d}t>\int_{0}^{1}\frac{t^{a-1}}{2}\,\mathrm{d}t=\frac{1}{2a},$ (31)
which leads to
$\frac{\partial\log\phi(a,b)}{\partial a}>0.$ (32)
Therefore, $\log\phi(a,b)$ increases monotonically in $a$ and so does $\phi(a,b)$. By symmetry, $\phi(a,b)$ also increases monotonically in $b$.
A.5 Proof of Proposition 1(iii)
To see the 1-1 mapping between $(r,\phi)\in(0,1)^{2}$ and $(a,b)\in(0,+\infty)^{2}$, it suffices to provide an injective and surjective mapping from $(a,b)$ to $(r,\phi)$. Define the mapping $h:(a,b)\to(r,\phi)$ as
$h:(a,b)\to\left(r=\frac{a}{a+b},\ \phi=\tilde{\phi}(a)\tilde{\phi}(b)\right),$ (33)
where $\tilde{\phi}(x)$ is defined in Appendix A.4.
Before showing that $h$ is injective and surjective, we first establish the properties of the restricted univariate mapping from $a$ to $\phi$ for arbitrarily fixed $r\in(0,1)$ induced by $h$, denoted as $\phi_{h}(a;r)$. For fixed $r$, we have b=a(1r)/rb=a(1-r)/r, and $h$ maps $(a,b)$ to $(r,\phi_{h}(a;r))$, where $\phi_{h}(a;r)$ is defined as
$\phi_{h}(a;r)=\tilde{\phi}(a)\tilde{\phi}(b)=\tilde{\phi}(a)\tilde{\phi}\left(a(1-r)/r\right).$ (34)
First, $\phi_{h}(a;r)$ increases monotonically in $a$, because
$\frac{d\log\phi_{h}(a;r)}{da}=\frac{d\log\tilde{\phi}(a)}{da}+\frac{d\log\tilde{\phi}(b)}{da}=\frac{d\log\tilde{\phi}(a)}{da}+\frac{d\log\tilde{\phi}(b)}{db}\frac{db}{da}>0,$ (35)
by Equation (29), Equation (32), and db/da=(1r)/r>0db/da=(1-r)/r>0.
Second, the image of $\phi_{h}(a;r)$ is $(0,1)$, i.e., $\phi_{h}(a;r)$ maps $a\in(0,+\infty)$ onto the domain of $\phi$ excluding 0 and $1$. Evaluate the limit of $\tilde{\phi}(x)$ at $x\to 0^{+}$ and $x\to+\infty$ respectively, we have
$\lim_{x\to 0^{+}}\tilde{\phi}(x)=\lim_{x\to 0^{+}}\frac{\Gamma(x+0.5)}{\sqrt{x}\Gamma(x)}=\lim_{x\to 0^{+}}\sqrt{x}\frac{\Gamma(x+0.5)}{\Gamma(x+1)}=0\times\frac{\sqrt{\pi}}{1}=0,$ (36)
by using $\Gamma(x)=\Gamma(x+1)/x$, and
$\lim_{x\to+\infty}\tilde{\phi}(x)=\lim_{x\to+\infty}\frac{\Gamma(x+0.5)}{\sqrt{x}\Gamma(x)}=1,$ (37)
by Equation (12) in Wendel (1948). Since both limits are finite and $a\to 0^{+}$ (or $\infty$) implies $b\to 0^{+}$ (or $\infty$) respectively, therefore
$\lim_{a\to 0^{+}}\phi_{h}(a;r)=0,\quad\lim_{a\to+\infty}\phi_{h}(a;r)=1.$ (38)
Given that $\phi_{h}(a;r)$ is continuous and monotonically increasing in $a$, by intermediate value theorem, for any $\phi^{*}\in(0,1)$ there exists a unique $a^{*}$ such that $\phi_{h}(a^{*};r)=\phi^{*}$. Therefore, the image of $\phi_{h}(a;r)$ is $(0,1)$.
Now, we show $h$ is injective and surjective. For injectivity, suppose that $h$ maps $(a_{1},b_{1})$ and $(a_{2},b_{2})$ to the same $(r,\phi)$. Therefore, $\phi_{h}(a_{1};r)=\phi_{h}(a_{2};r)=\phi$. Since $\phi_{h}(a;r)$ is monotonic in $a$, so $a_{1}=a_{2}$, hence $b_{1}=a_{1}(1-r)/r=a_{2}(1-r)/r=b_{2}$, establishing injectivity.
For surjectivity, pick an arbitrary $(r^{*},\phi^{*})\in(0,1)^{2}$. It induces $\phi_{h}(a;r^{*})$ which has the image of $(0,1)$, so there exists $a^{*}$ such that $\phi_{h}(a^{*};r^{*})=\phi^{*}$. Define $b^{*}=a^{*}(1-r^{*})/r^{*}$, then $h$ maps $(a^{*},b^{*})$ to $(r^{*},\phi^{*})$, establishing surjectivity.
A.6 Proof of Equation (12)
Recall $e(X)=\operatorname{\mathrm{e}xpit}(W_{e})$ and $W_{e}\sim\mathsf{N}(\mu_{e},\sigma_{e}^{2})$ given in Section 3.2, and $Y(z)=aW_{e}+z\tau+\epsilon$ given in Section 3.3, where $\epsilon$ is the compound error and $\operatorname{\mathbb{E}}[\epsilon]=\mu$ and $\operatorname{\mathbb{V}}[\epsilon]=\sigma_{y}^{2}$. For clarity of the proof, we write out $\epsilon=W_{e}^{\perp}+\epsilon_{ori}$, where $\epsilon_{ori}$ is the original error term defined in the RMM Model (11).
By the definition of Model (11), $\epsilon_{ori}$ is uncorrelated with any measurable function of $X$, including $W_{e}$, and therefore $\operatorname{\mathrm{C}ov}(W_{e},\epsilon)=\operatorname{\mathrm{C}ov}(W_{e},W_{e}^{\perp})+\operatorname{\mathrm{C}ov}(W_{e},\epsilon_{ori})=0$. Hence we have the following results:
$S^{2}$ $\equiv\mathbb{V}[Y(0)]=\mathbb{V}[aW_{e}+\epsilon]=a^{2}\sigma_{e}^{2}+\sigma_{y}^{2};$ (39)
$\rho$ $\equiv\mathrm{cor}[Y(0),W_{e}]=\mathrm{cor}[aW_{e}+\epsilon,W_{e}]=a\sqrt{\sigma_{e}^{2}/S^{2}}.$ (40)
By (40), $a=\rho\sqrt{{S^{2}}/\sigma_{e}^{2}}$, which, plugged into (39), yields
$\sigma_{y}^{2}=(1-\rho^{2})S^{2}.$ (41)
A.7 Proof of the analytical form of variance $V$ in Theorem 2
@@
$V=\mathbb{E}\left[\frac{\{Y(1)-\mathbb{E}[Y(1)]\}^{2}}{e(X)}\right]+\mathbb{E}\left[\frac{\{Y(0)-\mathbb{E}[Y(0)]\}^{2}}{1-e(X)}\right],$ (42)
where $e(X)$ and $Y(z)$ are given in Appendix A.6. We first state all necessary results regarding $\epsilon$, $\epsilon_{ori}$, $W_{e}$, and $W_{e}^{\perp}$ in Lemma 4.
Lemma 4.
Under the joint Gaussianity of $(f(X),W_{e})$ and the RMM Model 11:
  1. (1)
    $\operatorname{\mathbb{E}}[\epsilon_{ori}|W_{e}]=0$
  2. (2)
    $\mu=\operatorname{\mathbb{E}}[\epsilon]=\operatorname{\mathbb{E}}[\epsilon|W_{e}]=\operatorname{\mathbb{E}}[W_{e}^{\perp}]$
  3. (3)
    $\operatorname{\mathbb{E}}[\epsilon_{ori}|W_{e},W_{e}^{\perp}]=0$
  4. (4)
    $\operatorname{\mathrm{C}ov}(W_{e}^{\perp},\epsilon_{ori})=0$
Lemma 4 holds because: (1) $\operatorname{\mathbb{E}}[\epsilon_{ori}|X]=0$ and the law of iterated expectation, see Section 2.2.4 of Wooldridge (2010), that $\operatorname{\mathbb{E}}[\epsilon_{ori}|W_{e}]=\operatorname{\mathbb{E}}\left\{\operatorname{\mathbb{E}}[\epsilon_{ori}|X]\mid W_{e}\right\}=0$. (2) $W_{e}$ and $W_{e}^{\perp}$ are independent, and $\operatorname{\mathbb{E}}[\epsilon|W_{e}]=\operatorname{\mathbb{E}}[W_{e}^{\perp}|W_{e}]+\operatorname{\mathbb{E}}[\epsilon_{ori}|W_{e}]=\operatorname{\mathbb{E}}[W_{e}^{\perp}]$. (3) $\operatorname{\mathbb{E}}[\epsilon_{ori}|W_{e},W_{e}^{\perp}]=\operatorname{\mathbb{E}}[\epsilon_{ori}|W_{e},f(X)]=\operatorname{\mathbb{E}}\left\{\operatorname{\mathbb{E}}[\epsilon_{ori}|X]\mid W_{e},f(X)\right\}=0$, see that any one of $(f(X),W_{e},W_{e}^{\perp})$ is determined by another two. (4) $W_{e}^{\perp}=f(X)-aW_{e}$ is a function of $X$ and therefore uncorrelated with $\epsilon_{ori}$ by the definition of RMM.
We derive the first term in $V$, and similar result holds for the second term by symmetry.
@@
$\mathbb{E}\left[\frac{\{Y(1)-\mathbb{E}[Y(1)]\}^{2}}{e(X)}\right]$ $=\mathbb{E}\left[\frac{\{a(W_{e}-\mu_{e})+(\epsilon-\mu)\}^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ (43)
@@
$=a^{2}\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\mathbb{E}\left[\frac{(\epsilon-\mu)^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]+2a\mathbb{E}\left[\frac{(W_{e}-\mu_{e})(\epsilon-\mu)}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ (44)
$:=A+B+C$ (45)
Term $A$ only depends on the propensity score model but not the outcome model, while $B$ and $C$ involve both models. We first derive $B$ and $C$.
@@
$B$ $=\mathbb{E}\left\{\mathbb{E}\left[\frac{(\epsilon-\mu)^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\Bigg|W_{e}\right]\right\}=\mathbb{E}\left\{\frac{\operatorname{\mathbb{E}}[(\epsilon-\mu)^{2}\mid W_{e}]}{\operatorname{\mathrm{e}xpit}(W_{e})}\right\}=\mathbb{E}\left\{\frac{\operatorname{\mathbb{V}}[\epsilon\mid W_{e}]}{\operatorname{\mathrm{e}xpit}(W_{e})}\right\},$ (46)
and for $\operatorname{\mathbb{V}}[\epsilon\mid W_{e}]$ we have
$\operatorname{\mathbb{V}}[\epsilon\mid W_{e}]$ $=\operatorname{\mathbb{V}}[W_{e}^{\perp}|W_{e}]+\operatorname{\mathbb{V}}[\epsilon_{ori}|W_{e}]+2\operatorname{\mathrm{C}ov}(W_{e}^{\perp},\epsilon_{ori}|W_{e}).$ (47)
For these three pieces: $\operatorname{\mathbb{V}}[W_{e}^{\perp}|W_{e}]=\operatorname{\mathbb{V}}[W_{e}^{\perp}]$ because $W_{e}$ is independent of $W_{e}^{\perp}$; $\operatorname{\mathbb{V}}[\epsilon_{ori}|W_{e}]=\operatorname{\mathbb{V}}[\epsilon_{ori}]$ because of the homoskedasticity assumption that $\operatorname{\mathbb{V}}[\epsilon_{ori}|e(X)]=\operatorname{\mathbb{V}}[\epsilon_{ori}]$, and $W_{e}$ is a deterministic function of $e(X)$; and for the last piece,
$\operatorname{\mathrm{C}ov}(W_{e}^{\perp},\epsilon_{ori}|W_{e})$ $=\operatorname{\mathbb{E}}[W_{e}^{\perp}\cdot\epsilon_{ori}|W_{e}]-\operatorname{\mathbb{E}}[W_{e}^{\perp}|W_{e}]\operatorname{\mathbb{E}}[\epsilon_{ori}|W_{e}]$ (48)
[…]
Therefore, $\operatorname{\mathbb{V}}[\epsilon\mid W_{e}]$ is invariant to $W_{e}$ by seeing that
$\operatorname{\mathbb{V}}[\epsilon\mid W_{e}]=\operatorname{\mathbb{V}}[W_{e}^{\perp}]+\operatorname{\mathbb{V}}[\epsilon_{ori}]+0=\operatorname{\mathbb{V}}[\epsilon]=\sigma_{y}^{2},$ (51)
and consequently
@@
$B=\mathbb{E}\left\{\frac{\operatorname{\mathbb{V}}[\epsilon\mid W_{e}]}{\operatorname{\mathrm{e}xpit}(W_{e})}\right\}=\sigma_{y}^{2}\cdot\mathbb{E}\left[\frac{1}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ (52)
Finally, for $C$,
$C$ $=2a\cdot\mathbb{E}\left\{\frac{W_{e}-\mu_{e}}{\operatorname{\mathrm{e}xpit}(W_{e})}\operatorname{\mathbb{E}}[\epsilon-\mu\mid W_{e}]\right\}$ (53)
[…]
@@
$=2a\cdot\mathbb{E}\left\{\frac{W_{e}-\mu_{e}}{\operatorname{\mathrm{e}xpit}(W_{e})}\cdot 0\right\}=0.$ (55)
Gathering $A$, $B$, and $C$ above,
@@
$\mathbb{E}\left[\frac{\{Y(1)-\mathbb{E}[Y(1)]\}^{2}}{e(X)}\right]=a^{2}\cdot\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{y}^{2}\cdot\mathbb{E}\left[\frac{1}{\operatorname{\mathrm{e}xpit}(W_{e})}\right],$ (56)
which is fully determined under $a^{2}$, $\sigma_{y}^{2}$, and the distribution of $W_{e}$, as calculated below.
$\mathbb{E}\left[\frac{1}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ $=\int\left(1+e^{-w}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(w-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}w$ (57)
$=\int\left(1+e^{-u-\mu_{e}}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u$ (58)
$=1+\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right),$ (59)
$\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ $=\int(w-\mu_{e})^{2}\left(1+e^{-w}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(w-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}w$ (60)
$=\int u^{2}\left(1+e^{-u-\mu_{e}}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u$ (61)
$=\sigma_{e}^{2}+(\sigma_{e}^{4}+\sigma_{e}^{2})\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (62)
@@
$\mathbb{E}\left[\frac{\{Y(1)-\mathbb{E}[Y(1)]\}^{2}}{e(X)}\right]=a^{2}\sigma_{e}^{2}+\sigma_{y}^{2}+\left[a^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{y}^{2}\right]\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (63)